# Revisiting wideband pulsar timing measurements

### v2

@@ Abstract
pubyear: 2025pagerange: Revisiting wideband pulsar timing measurementsC
@@ 1 Introduction
Pulsars are rotating neutron stars whose electromagnetic radiation appears as periodic pulses to a terrestrial observer (LorimerKramer2012). They are characterized by high rotational stability, acting as accurate celestial clocks. Pulsar timing is the technique of tracking a pulsar’s rotation by measuring the times of arrival (TOAs) of its pulses (HobbsEdwardsManchester2006). Pulsar timing is one of the most precise techniques in astronomy with a wide range of applications, from spacecraft navigation (DengHobbs+2013) to gravitational wave detection (AgazieAntoniadis+2024).
@@
Radio signals emitted by a pulsar are dispersed as they propagate through the ionized interstellar medium (ISM), introducing a delay that is proportional to the electron column density along the line of sight and inversely proportional to the square of the radio frequency (BackerHellings1986). The electron column density towards a pulsar, known as the dispersion measure (DM), varies stochastically as a function of time due to the dynamic nature of the ISM and the relative motion of the pulsar with respect to the Earth. In addition to the pulsar rotation and interstellar dispersion, the TOAs measured at a terrestrial observatory are influenced by several astrophysical phenomena, including geometric and relativistic delays due to the orbital motion of the pulsar and the Earth, passing gravitational waves, and instrumental delays and noise (HobbsEdwardsManchester2006; EdwardsHobbsManchester2006; DamourDeruelle1986; EstabrookWahlquist1975). Pulsar timing involves characterizing all of these effects to construct a predictive model for the TOAs.
@@
A pulsar observation using a radio telescope generates an observing frequency-resolved radio light curve. This light curve is folded using the known rotational frequency of the pulsar to obtain an integrated pulse profile, which gives the pulsar intensity as a function of the pulse phase and the observing frequency. In conventional pulsar timing, also known as narrowband timing, a frequency-resolved integrated pulse profile spanning multiple frequency sub-bands covering the observed band-pass is produced, and a TOA is measured from each sub-band independently by cross-correlating it against a similarly frequency-resolved template profile (Taylor1992). On the other hand, in the wideband timing paradigm, a single TOA along with a DM is measured from a frequency-resolved two-dimensional integrated pulse profile, also known as a portrait, by cross-correlating it against a two-dimensional template (PennucciDemorestRansom2014; LiuDesvignes+2014). Wideband timing can also be performed on combined datasets consisting of simultaneous multi-band observations (PaladiDwivedi+2024). The two-dimensional templates for wideband timing, also known as template portraits, are usually obtained using a principal component analysis-based method, described in Pennucci2019. In general, a wideband timing measurement produces the TOA and DM estimates, their uncertainties, and their covariance, although it turns out that the measurement covariance can be made to vanish by adjusting the fiducial frequency at which the measurement is made; see Section 2. Consequently, wideband timing and noise analysis are done by incorporating both the TOA and the DM measurements (AlamArzoumanian+2021; SusobhananVan_Haasteren2025; Susobhanan2025b). Wideband timing provides two main advantages over narrowband timing — reduced data volumes and better handling of frequency-dependent profile variability (HankinsRickett1986).
@@
In this work, we present an improved method for wideband TOA and DM measurements, which provides better handling of the noise present in the observed portrait, thereby producing more realistic estimates of the measurement uncertainty. This paper is arranged as follows. In Section 2, we derive our new method for wideband measurements, starting from first principles. In Section 3, we demonstrate our method using a simulated example profile. In Section 4, we demonstrate our methods using data obtained as part of the Indian Pulsar Timing Array (InPTA) experiment (JoshiArumugasamy+2018; TarafdarNobleson+2022; RanaTarafdar+2025). Finally, we summarize our results in Section 5.
@@ 2 Revisiting wideband measurements
We begin by writing down a model for the observed total intensity portrait as a function of the pulse phase $\phi$ and the observing frequency $\nu$ following PennucciDemorestRansom2014.
@@
where $P(\phi,\nu)$ is the observed portrait, $T(\phi,\nu)$ is the template portrait, $b(\nu)$ is the baseline total intensity, $a(\nu)$ is the amplitude of the signal at each frequency, $\varphi(\nu)$ is a frequency-dependent phase shift, and $n(\phi,\nu)$ is a white noise process. The frequency dependence of $a(\nu)$ and $b(\nu)$ arises because of the pulsar’s intrinsic emission spectrum (MaronKijak+2000), interstellar scintillation (Rickett1970), and the shape of the telescope bandpass. The phase shift $\varphi(\nu)$ is given by
@@
In practice, the observed portrait is discrete, and equation (1) can be written in a discrete form as
@@
where $\alpha\in\{1,...,N_{\text{chan}}\}$ represents the observing frequency channel, $N_{\text{chan}}$ is the number of frequency channels, $j\in\{1,...,N_{\text{bin}}\}$ represents the pulse phase bin, $N_{\text{bin}}$ is the number of phase bins, and $\varphi_{\alpha}=\varphi_{0}-\kappa D\bar{F}\left(\nu_{\alpha}^{-2}-\nu_{\text{ref}}^{-2}\right)$. Following Taylor1992 and PennucciDemorestRansom2014, we compute the discrete Fourier transform of equation (4) along the phase index. Applying the Fourier shift theorem and discarding the zero-frequency baseline term, we obtain
@@
$\ln L$ is a function of the parameters of interest $\varphi_{0}$ and $D$, the nuisance parameters $a_{\alpha}$ and $\sigma_{\alpha}$, and the arbitrary constant $\nu_{\text{ref}}$. PennucciDemorestRansom2014 removes the $a_{\alpha}$ dependence by maximizing equation (6) over those parameters, and fixes $\nu_{\text{ref}}$ by demanding that the measurement covariance between $\varphi_{0}$ and $D$ vanish. Multiple methods have been employed to estimate the noise standard deviation $\sigma_{\alpha}$. This includes using the variance within a window around the profile’s minimum, assuming it lies in the off-pulse region, or estimating the variance based on the highest harmonics in the discrete Fourier transform $\tilde{P}_{\alpha k}$ (PennucciDemorestRansom2014). The latter method is implemented in PulsePortraiture. Unfortunately, these approaches do not always provide reliable noise estimates: for example, the former method fails for pulsars with large duty cycles and the latter fails when the profile signal-to-noise ratio is sufficiently high, making even the higher harmonics dominated by the signal. The latter failure mode occurs when (a) the pulsar is bright, (b) the telescope has a large collecting area, and/or (c) the pulsar is observed for a long duration.
@@
where $\mathfrak{p}$ denotes the posterior distribution, the likelihood $L$ is given by equation (6), $\Pi$ denotes the prior distributions, ZZ is the Bayesian evidence, and $\boldsymbol{a}$ and $\boldsymbol{\sigma}$ represent the collection of $a_{\alpha}$ and $\sigma_{\alpha}$ parameters respectively. We can marginalize equation (8) analytically over $a_{\alpha}$ and $\sigma_{\alpha}$ by imposing the appropriate conjugate priors on those parameters. It turns out, from equation (6), that the conjugate prior on $a_{\alpha}$ is a Gaussian distribution and the conjugate prior on $\sigma_{\alpha}$ is an inverse-gamma distribution, whose probability density function is given by
$\Pi(\sigma_{\alpha};p,q)=\frac{q^{p}}{\Gamma(p)}\sigma_{\alpha}^{-p-1}\exp(-\sigma_{\alpha}/x)\,.$
Since we do not have any prior information on $a_{\alpha}$ or $\sigma_{\alpha}$ except that the latter should be positive, we choose $\Pi[a_{\alpha}]$ to be a Gaussian distribution with infinite variance (same as an infinitely wide uniform distribution), i.e., $\Pi[a_{\alpha}]\propto 1$. Further, we choose an improper log-uniform prior for $\sigma_{\alpha}$, corresponding to $\Pi[\sigma_{\alpha}]\propto\sigma_{\alpha}^{-1}$, which is the same as the limiting case of an inverse-gamma distribution with $p,q\rightarrow 0$ (this is the same as the Jeffreys’ prior; Jeffreys1946).
Marginalizing equation (8) over the $a_{\alpha}$ and $\sigma_{\alpha}$ using the above-mentioned prior distributions, we obtain
@@
We note in passing that an alternative to marginalization is the profile likelihood approach (HeroldFerreiraHeinrich2025), where the likelihood function is maximized over the nuisance parameters. This approach yields an expression that is very similar to equation (10):
@@
Since the fiducial frequency $\nu_{\text{ref}}$ is arbitrary, it is usually chosen such that the measurement covariance between $\varphi_{0}$ and $D$ vanishes, so that they can be treated as independent measurements, in turn simplifying and reducing the computational cost of wideband timing and noise analysis (AlamArzoumanian+2021; SusobhananVan_Haasteren2025). This is achieved in PennucciDemorestRansom2014 using a Hessian-based analytic estimate of the measurement covariance, by setting the second derivative $\frac{\partial^{2}\ln L}{\partial\varphi_{0}\partial D}$ to zero. This choice of the fiducial frequency for our noise-marginalized likelihood, denoted $\bar{\nu}_{\text{ref}}$, is derived in Appendix C. It can be seen from Appendix C that (a) the $\bar{\nu}_{\text{ref}}$ depends on the observed portrait on each epoch, and therefore is different for each epoch, and (b) the $\bar{\nu}_{\text{ref}}$ estimate is derived assuming a Gaussian posterior distribution, and may not be accurate when the posterior is significantly non-Gaussian, e.g., in the presence of loud RFI. The measurement uncertainties in $\varphi_{0}$ and $D$ can also be estimated with the help of the Hessian, and these computations are also described in Appendix C.
@@ 3 A simulated example
Figure 1: The top plot shows the simulated pulse portrait based on the analytic two-peaked template given in equation (12), also containing white noise in each frequency channel. Ten randomly selected channels have noise injections with a significantly higher amplitude, simulating RFI, and these are visible in the plot as horizontal stripes. The bottom plot shows the post-fit residuals of the portrait, which are determined by subtracting the best-fitting model from the simulated data. The residuals show only noise, indicating a successful fit.
We now turn to an example to showcase our method using a simulated portrait. We assume a frequency-dependent two-peaked template portrait of the form
@@
We simulate an observed portrait using equation (4) with $b(\nu)=0.2$, $a(\nu)=2$, $\varphi_{0}=0.1$, $D=0.15$ pc/cm3, $\nu_{\text{ref}}=750$ MHz, $\bar{F}=100$ Hz, $N_{\text{bin}}=128$, and $N_{\text{chan}}=256$ for observing frequencies between 500–1000 MHz. We inject white noise realizations with a standard deviation $\sigma_{\alpha}=0.025$ in every frequency channel, except for 10 randomly selected channels where we use $\sigma_{\alpha}=0.25$ to simulate RFI. The resulting simulated portrait is shown in Figure 1.
@@
We estimate $\varphi_{0}$ and $D$ from this simulated template using the amplitude and noise-marginalized likelihood function of equation (10). Assuming broad uniform priors for $\varphi_{0}$ and $D$, we draw samples from the posterior distribution (equation 9) using the MultiNest nested sampling algorithm (FerozHobsonBridges2009), implemented in the nestle package (Barbary2021). The posterior distribution $p[\varphi_{0},D|P,T]$ at a fiducial frequency $\bar{\nu}_{\text{ref}}$ estimated based on Appendix C is plotted in Figure 2. The portrait residuals, obtained by subtracting the best-fit model from the data portrait in this case, are plotted in the lower panel of Figure 1. We find that the injected $\varphi_{0}$ and $D$ values are reliably estimated, and that the residuals in each frequency channel are white.
@@
We now proceed to ascertain the accuracy of the $\bar{\nu}_{\text{ref}}$ obtained in Appendix C. For this purpose, we estimate $\varphi_{0}$ and $D$ using 101 uniformly separated values of $\nu_{\text{ref}}$ in the 400–1100 MHz range using the noise-marginalized likelihood of equation (10) using nested sampling. We find that the estimated $\phi_{0}$ and $D$ values are consistent with the injected values for every value of $\nu_{\text{ref}}$, as expected [^1] The measured $\phi_{0}$ value depends on $\nu_{\text{ref}}$. We deem the measured $\phi_{0}$ value to be consistent with the injected value if $\left|\phi_{0}^{\text{meas}}-\phi_{0}^{\text{inj}}-\kappa DF({\nu}_{\text{ref}}^{\text{meas}}-{\nu}_{\text{ref}}^{\text{inj}})\right|$ is less than the measurement uncertainty of $\phi_{0}$, where the superscripts ‘meas’ and ‘inj’ indicate the measured and injected values. On the other hand, the measured $D$ value should be independent of $\nu_{\text{ref}}$.. We plot the Pearson correlation coefficient between $\varphi_{0}$ and $D$ as a function of $\nu_{\text{ref}}$ in Figure 3. We find that the correlation becomes approximately zero at $\nu_{\text{ref}}=\bar{\nu}_{\text{ref}}$, although a tiny correlation remains even at this $\nu_{\text{ref}}$ value [^2] A similar observation was also made by PennucciDemorestRansom2014. We also plot the determinant of the covariance matrix, which is a measure of the total uncertainty in the wideband measurement, as a function of $\nu_{\text{ref}}$, in Figure 3. We find that this determinant value remains independent of $\nu_{\text{ref}}$ up to numerical precision, indicating that the overall measurement precision is not affected by the choice of $\nu_{\text{ref}}$.

### v3

@@ Abstract
pubyear: 2026pagerange: Revisiting wideband pulsar timing measurementsC
@@ 1 Introduction
Pulsars are rotating neutron stars whose electromagnetic radiation appears as periodic pulses to a terrestrial observer (Lorimer and Kramer, 2012). They are characterized by high rotational stability, acting as accurate celestial clocks. Pulsar timing is the technique of tracking a pulsar’s rotation by measuring the times of arrival (TOAs) of its pulses (Hobbs et al., 2006). Pulsar timing is one of the most precise techniques in astronomy with a wide range of applications, from spacecraft navigation (Deng et al., 2013) to gravitational wave detection (Agazie et al., 2024).
@@
Radio signals emitted by a pulsar are dispersed as they propagate through the ionized interstellar medium (ISM), introducing a delay that is proportional to the electron column density along the line of sight and inversely proportional to the square of the radio frequency (Backer and Hellings, 1986). The electron column density towards a pulsar, known as the dispersion measure (DM), varies stochastically as a function of time due to the dynamic nature of the ISM and the relative motion of the pulsar with respect to the Earth. In addition to the pulsar rotation and interstellar dispersion, the TOAs measured at a terrestrial observatory are influenced by several astrophysical phenomena, including geometric and relativistic delays due to the orbital motion of the pulsar and the Earth, passing gravitational waves, and instrumental delays and noise (Hobbs et al., 2006; Edwards et al., 2006; Damour and Deruelle, 1986; Estabrook and Wahlquist, 1975). Pulsar timing involves characterizing all of these effects to construct a predictive model for the TOAs.
@@
A pulsar observation using a radio telescope generates an observing frequency-resolved radio light curve. This light curve is folded using the known rotational frequency of the pulsar to obtain an integrated pulse profile, which gives the pulsar intensity as a function of the pulse phase and the observing frequency. In conventional pulsar timing, also known as narrowband timing, a frequency-resolved integrated pulse profile spanning multiple frequency sub-bands covering the observed band-pass is produced, and a TOA is measured from each sub-band independently by cross-correlating it against a similarly frequency-resolved template profile (Taylor, 1992). On the other hand, in the wideband timing paradigm, a single TOA along with a DM is measured from a frequency-resolved two-dimensional integrated pulse profile, also known as a portrait, by cross-correlating it against a two-dimensional template (Pennucci et al., 2014; Liu et al., 2014). Wideband timing can also be performed on combined datasets consisting of simultaneous multi-band observations (Paladi et al., 2024). The two-dimensional templates for wideband timing, also known as template portraits, are usually obtained using a principal component analysis-based method, described in Pennucci (2019). In general, a wideband timing measurement produces the TOA and DM estimates, their uncertainties, and their covariance, although it turns out that the measurement covariance can be made to vanish by adjusting the fiducial frequency at which the measurement is made; see Section 2. Consequently, wideband timing and noise analysis are done by incorporating both the TOA and the DM measurements (Alam et al., 2021; Susobhanan and van Haasteren, 2025; Susobhanan, 2025). Algorithms for generating template portraits and for performing wideband TOA and DM measurements are implemented in the PulsePortraiture package (Pennucci et al., 2014; Pennucci, 2019), wideband timing methods are implemented in tempo (Nice et al., 2015) and PINT (Luo et al., 2021; Susobhanan et al., 2024), and Bayesian wideband noise analysis is implemented in ENTERPRISE (Ellis et al., 2020) and Vela.jl (Susobhanan, 2025). Wideband timing provides two main advantages over narrowband timing — reduced data volumes and better handling of frequency-dependent profile variability (Hankins and Rickett, 1986).
@@
In this work, we present an improved method for wideband TOA and DM measurements, which provides better handling of the noise present in the observed portrait, thereby producing more realistic estimates of the measurement uncertainty. This paper is arranged as follows. In Section 2, we derive our new method for wideband measurements, starting from first principles. In Section 3, we demonstrate our method using a simulated example profile. In Section 4, we demonstrate our methods using data obtained as part of the Indian Pulsar Timing Array (InPTA) experiment (Joshi et al., 2018; Tarafdar et al., 2022; Rana et al., 2025). Finally, we summarize our results in Section 5.
@@ 2 Revisiting wideband measurements
We begin by writing down a model for the observed total intensity portrait as a function of the pulse phase $\phi$ and the observing frequency $\nu$ following Pennucci et al. (2014).
@@
where $P(\phi,\nu)$ is the observed portrait, $T(\phi,\nu)$ is the template portrait, $b(\nu)$ is the baseline total intensity, $a(\nu)$ is the amplitude of the signal at each frequency, $\varphi(\nu)$ is a frequency-dependent phase shift, and $n(\phi,\nu)$ is a white noise process. The frequency dependence of $a(\nu)$ and $b(\nu)$ arises because of the pulsar’s intrinsic emission spectrum (Maron et al., 2000), interstellar scintillation (Rickett, 1970), and the shape of the telescope bandpass. The phase shift $\varphi(\nu)$ is given by
@@
In practice, the observed portrait is discrete, and equation (1) can be written in discrete form as
@@
where $\alpha\in\{1,...,N_{\text{chan}}\}$ represents the observing frequency channel, $N_{\text{chan}}$ is the number of frequency channels, $j\in\{1,...,N_{\text{bin}}\}$ represents the pulse phase bin, $N_{\text{bin}}$ is the number of phase bins, and $\varphi_{\alpha}=\varphi_{0}-\kappa D\bar{F}\left(\nu_{\alpha}^{-2}-\nu_{\text{ref}}^{-2}\right)$. Following Taylor (1992) and Pennucci et al. (2014), we compute the discrete Fourier transform of equation (4) along the phase index. Applying the Fourier shift theorem and discarding the zero-frequency baseline term, we obtain
@@
$\ln L$ is a function of the parameters of interest $\varphi_{0}$ and $D$, the nuisance parameters $a_{\alpha}$ and $\sigma_{\alpha}$, and the arbitrary constant $\nu_{\text{ref}}$. Pennucci et al. (2014) removes the $a_{\alpha}$ dependence by maximizing equation (6) over those parameters, and fixes $\nu_{\text{ref}}$ by demanding that the measurement covariance between $\varphi_{0}$ and $D$ vanish. Multiple methods have been employed to estimate the noise standard deviation $\sigma_{\alpha}$. This includes using the variance within a window around the profile’s minimum, assuming it lies in the off-pulse region, or estimating the variance based on the highest harmonics in the discrete Fourier transform $\tilde{P}_{\alpha k}$ (Pennucci et al., 2014). The latter method is implemented in PulsePortraiture. Unfortunately, these approaches do not always provide reliable noise estimates: for example, the former method fails for pulsars with large duty cycles and the latter fails when the profile signal-to-noise ratio is sufficiently high, making even the higher harmonics dominated by the signal. The latter failure mode occurs when (a) the pulsar is bright, (b) the telescope has a large collecting area, and/or (c) the pulsar is observed for a long duration.
@@
where $\mathfrak{p}$ denotes the posterior distribution, the likelihood $L$ is given by equation (6), $\Pi$ denotes the prior distributions, ZZ is the Bayesian evidence, and $\boldsymbol{a}$ and $\boldsymbol{\sigma}$ represent the collection of $a_{\alpha}$ and $\sigma_{\alpha}$ parameters respectively. We can marginalize equation (8) analytically over $a_{\alpha}$ and $\sigma_{\alpha}$ by imposing the appropriate conjugate priors on those parameters. We choose the Jeffreys’ priors (Jeffreys, 1946) for these parameters, which correspond to an improper uniform prior for $a_{\alpha}$ and an improper log-uniform prior for $\sigma_{\alpha}$; i.e.,
$\Pi[a_{\alpha}]$ $\propto 1\,$, (9a)
$\Pi[\sigma_{\alpha}]$ $\propto\sigma_{\alpha}^{-1}\,.$ (9b)
Marginalizing equation (8) over the $a_{\alpha}$ and $\sigma_{\alpha}$ using the above-mentioned priors, we obtain
@@
We note in passing that an alternative to marginalization is the profile likelihood approach (Herold et al., 2025), where the likelihood function is maximized over the nuisance parameters. This approach yields an expression that is very similar to equation (11):
@@
Since the fiducial frequency $\nu_{\text{ref}}$ is arbitrary, it is usually chosen such that the measurement covariance between $\varphi_{0}$ and $D$ vanishes, so that they can be treated as independent measurements, in turn simplifying and reducing the computational cost of wideband timing and noise analysis (Alam et al., 2021; Susobhanan and van Haasteren, 2025). This is achieved in Pennucci et al. (2014) using a Hessian-based analytic estimate of the measurement covariance, by setting the second derivative $\frac{\partial^{2}\ln L}{\partial\varphi_{0}\partial D}$ to zero. This choice of the fiducial frequency for our noise-marginalized likelihood, denoted $\bar{\nu}_{\text{ref}}$, is derived in Appendix C. It can be seen from Appendix C that (a) the $\bar{\nu}_{\text{ref}}$ depends on the observed portrait on each epoch, and therefore is different for each epoch, and (b) the $\bar{\nu}_{\text{ref}}$ estimate is derived assuming a Gaussian posterior distribution, and may not be accurate when the posterior is significantly non-Gaussian, e.g., in the presence of loud RFI. The measurement uncertainties in $\varphi_{0}$ and $D$ can also be estimated with the help of the Hessian, and these computations are also given in Appendix C.
@@ 3 A simulated example
Figure 1: The top plot shows the simulated pulse portrait based on the analytic two-peaked template given in equation (13). The template is weighted by a dynamic spectrum emulating scintillation. It also contains white noise in each frequency channel. Ten randomly selected channels have noise injections with a significantly higher amplitude, simulating RFI, and these are visible in the plot as horizontal stripes. The bottom plot shows the post-fit residuals of the portrait, which are determined by subtracting the best-fitting model from the simulated data. The residuals show only noise, indicating a successful fit.
We now present an example to illustrate our method using a simulated portrait. We assume a frequency-dependent two-peaked template portrait of the form
@@
We simulate an observed portrait using equation (4) with $b(\nu)=0.2$, $\varphi_{0}=0.1$, $D=0.15$ pc/cm3, $\nu_{\text{ref}}=750$ MHz, $\bar{F}=100$ Hz, $N_{\text{bin}}=128$, and $N_{\text{chan}}=258$ for observing frequencies between 500–1000 MHz. We simulate frequency-dependent amplitude modulations due to interstellar scintillation using the SCINTOOLS package (Reardon et al., 2020) [^1] We set the Max Born scattering strength parameter ($m_{b}^{2}$) to be 15 and the structure function exponent to be 5/3.. We inject white noise realizations with a standard deviation $\sigma_{\alpha}=0.025$ in every frequency channel, except for 10 randomly selected channels where we use $\sigma_{\alpha}=0.25$ to simulate RFI. The resulting simulated portrait is shown in the top panel of Figure 1.
@@
We estimate $\varphi_{0}$ and $D$ from this simulated template using the amplitude and noise-marginalized likelihood function given in equation (11). Assuming broad uniform priors for $\varphi_{0}$ and $D$, we draw samples from the posterior distribution (equation 10) using the MultiNest nested sampling algorithm (Feroz et al., 2009), implemented in the nestle package (Barbary, 2021). The posterior distribution $p[\varphi_{0},D|P,T]$ at a fiducial frequency $\bar{\nu}_{\text{ref}}$ estimated based on Appendix C is plotted in Figure 2. The portrait residuals, obtained by subtracting the best-fit model from the data portrait in this case, are plotted in the bottom panel of Figure 1. We find that the injected $\varphi_{0}$ and $D$ values are reliably estimated, and that the residuals in each frequency channel are white.
@@
We now proceed to ascertain the accuracy of the $\bar{\nu}_{\text{ref}}$ obtained in Appendix C. For this purpose, we estimate $\varphi_{0}$ and $D$ using 101 uniformly separated values of $\nu_{\text{ref}}$ in the 400–1100 MHz range using the noise-marginalized likelihood of equation (11) using nested sampling. We find that the estimated $\phi_{0}$ and $D$ values are consistent with the injected values for every value of $\nu_{\text{ref}}$, as expected [^2] The measured $\phi_{0}$ value depends on $\nu_{\text{ref}}$. We deem the measured $\phi_{0}$ value to be consistent with the injected value if $\left|\phi_{0}^{\text{meas}}-\phi_{0}^{\text{inj}}-\kappa DF({\nu}_{\text{ref}}^{\text{meas}}-{\nu}_{\text{ref}}^{\text{inj}})\right|$ is less than the measurement uncertainty of $\phi_{0}$, where the superscripts ‘meas’ and ‘inj’ indicate the measured and injected values. On the other hand, the measured $D$ value should be independent of $\nu_{\text{ref}}$.. We plot the Pearson correlation coefficient between $\varphi_{0}$ and $D$ as a function of $\nu_{\text{ref}}$ in Figure 3. We find that the correlation becomes approximately zero at $\nu_{\text{ref}}=\bar{\nu}_{\text{ref}}$, although a tiny correlation remains even at this $\nu_{\text{ref}}$ value [^3] A similar observation was also made by Pennucci et al. (2014). We also plot the determinant of the covariance matrix, which is a measure of the total uncertainty in the wideband measurement, as a function of $\nu_{\text{ref}}$, in Figure 3. We find that this determinant value remains independent of $\nu_{\text{ref}}$ up to numerical precision, indicating that the overall measurement precision is not affected by the choice of $\nu_{\text{ref}}$.